[troffman,twoside]article document xvlFeb 26, 1992Rev: 2.11 NAME xv – interactive image display for the X Window System SYNTAX xv [options] [filename [filename…]] NOTE This man page is merely the text portion of the (PostScript) xv docs, without the figures,

and without the appendices. As such, it is highly recommended that you get your hands on the real documentation.

DESCRIPTION xv is an X11 program that displays images in the GIF, JPEG, TIFF, PBM, PGM, PPM, X11 bitmap,

PDS/VICAR, Sun Rasterfile, and PM formats on 1–, 2–, 4–, 6–, 8–, 16–, 24–, and 32–bit X displays. xv will also read compress-ed versions of these files.

SECTION 1: OVERVIEW xv version 2.11 lets you do a large number of things (many of them actually useful), including, but not

limited to, the following: TPlistbullet [ •] display an image in a window on the screen [ •] display an image on the root window, in a variety of styles [ •] arbitrarily stretch or compress the image [ •] rotate the image in 90-degree steps [ •] flip the image around the horizontal or vertical axes [ •] crop a rectangular portion of the image [ •] magnify any portion of the image by any amount, up to the size of the screen [ •] determine pixel values and x,y coordinates in the image [ •] adjust image brightness and contrast with a gamma correction function [ •] apply different gamma functions to the Red, Green, and Blue color components, to correct for nonlinear color response [ •] adjust global image saturation [ •] perform global hue remapping [ •] edit an image’s colormap [ •] reduce the number of colors in an image [ •] dither in color and b&w [ •] smooth an image [ •] crop off solid borders automatically [ •] convert image formats [ •] generate Encapsulated PostScript